Mr. Obama borrows money at 6% compound monthly . Two years later, he pays off the loan with a single payment of $5000. How much did he borrow?
Answer by stanbon(57984)

Mr. Obama borrows money at 6% compound monthly . Two years later, he pays off the loan with a single payment of $5000. How much did he borrow?
-------------
A(t) = P(1+(r/n))^(nt)
5000 = P(1+0.06/12)^(12*2
---
5000 = P(1.005)^24
P = 5000/1.12716
P = $4435.93 is the amount he borrowed.
=========================================
Cheers,
Stan H.